### v1

# Antichain cutsets in real-ranked lattices

Abstract.
We show that in a rank supersolvable lattice that is graded by a bounded real interval, any antichain cutset is a level set for some appropriately constructed grading. As a consequence, given an antichain cutset in any of the measurable Boolean lattice, a continuous partition lattice, or a continuous projective geometry, we may find a grading in which the cutset is a level set.
Work of the second author is supported in part by the Slovenian Research Agency (research program P1-0285 and research projects J1-9108, N1-0160, J1-2451).
1. Introduction
The Boolean lattice on a finite set is a basic example in lattice theory, with almost any nice property that one could ask for. The Boolean lattice on an infinite set, on the other hand, is much less well-behaved. For example, it is difficult or impossible to extend gradedness to this situation; in particular, ordinal theory gives that there are many isomorphism classes of maximal chains. Better-behaved is the measurable Boolean lattice on a complete measure space $(X,\mu)$ with a finite measure. In this lattice, we restrict ourselves to the measurable subsets of $X$, and consider two measurable subsets to be equivalent if their symmetric difference has measure zero. If the measure does not have any points of positive measure (or, in the other extreme, if $\mu$ is the counting measure), then $\mu$ gives an analogue of a rank function.
The situation with the partition lattice on an infinite set is similar: the full lattice of partitions on $[0,1]$ is too large to admit a nice rank function, and the same holds even for the lattice of partitions into measurable parts. But Von Neumann in [19, 20] developed machinery to build continuous extensions of lattices of subspaces and similar families of modular lattices. Indeed, the measurable Boolean lattice on the interval $[0,1]$ can be viewed as an example of Von Neumann’s construction. Björner in [2] extended the machinery of Von Neumann to geometric and semimodular lattices. A main application of Björner’s extension was a continuous extension of finite partition lattices. Haiman in [12] gave a more concrete construction of a continuous partition lattice, built from measurable sets in a similar manner to the measurable Boolean lattice. Björner’s lattice embeds in that of Haiman.
The current paper came out of an effort to extend some of our earlier work on antichain cutsets for discrete posets [8] to the continuous situation, covering (at least) all of the examples we have so far discussed. An antichain cutset in a lattice $L$ is an antichain $A$ so that every maximal chain nontrivially intersects $A$. Antichain cutsets were first studied by Rival and Zaguia [15]. The structure of antichain cutsets is much more prescribed than that of antichains more generally. For example, the family of antichain cutsets of a graded, finite, and strongly connected poset coincides exactly with that of the level sets of the poset [8, Theorem 2]. We contrast with the situation for general antichains, where e.g. a finite partition lattice typically has antichains that are larger than any level set [11].
Unlike the finite situation, where a rank function may be recovered in a straightforward way from the poset structure, the continuous lattices that we are interested in may admit gradings that are quite different from one another. For example, let $L$ be the measurable Boolean lattice on $[0,1]$ with Lebesgue measure $\mu$. Then $\mu$ is a rank function (see Section 2 for a definition), and given a measurable positive function $f$, the function $\nu(U)=\int_{U}f\,d\mu$ yields another rank function. The level sets of these two rank functions are generally quite different from each other. Since a level set of any rank function is necessarily an antichain cutset, it is too much to hope that every antichain cutset is a level set in some chosen rank function.
We believe that the following question is reasonable to ask, and of some interest.
Question 1.1.
For which classes of graded posets does it hold that, for each antichain cutset $A$, there is a grading $\sigma$ (possibly distinct from an initially presented one) so that $A$ is a level set under $\sigma$?
The first author showed in [7] that the condition asked for in Question 1.1 holds for space-time causality orders.
In the current paper, we will show that the condition of this question holds for a reasonably general class of lattices, defined as follows. Let $L$ be a lattice with an understood rank function $\rho:L\to R$, for $R$ some subset of $\mathbb{R}$, or more generally of the extended reals $\infty$$\infty$]. We say that an element $m$ of $L$ is rank modular if for every element $x$ of the lattice, it holds that $\rho(x\vee m)+\rho(x\wedge m)=\rho(x)+\rho(m)$. Equivalently, the same relation holds whenever $x$ is incomparable with $m$. We say that $L$ is rank supersolvable if it has a maximal chain consisting of rank modular elements (a chief chain).
In the case of a finite lattice with the natural rank function, rank supersolvability is one of several equivalent conditions for supersolvability [9]. We do not know if these conditions are equivalent in the non-discrete situation; see also Lemma 3.14 and the subsequent discussion.
Our main result is as follows.
Theorem 1.2.
Let $R\subseteq\mathbb{R}$ be a bounded interval. If $L$ is a rank supersolvable lattice with respect to the $R$-grading $\rho$, and $A$ is an antichain cutset for $L$, then there is some grading $\sigma$ under which $A$ is a level set.
Under an additional continuity condition, we can extend Theorem 1.2 to unbounded intervals. A precise statement of this extension may be found in Theorem 4.1 below.
Corollary 1.3.
For any antichain cutset $A$ in
  1. (1)
    the measurable Boolean lattice over $[0,1]$ with the Lesbesgue measure, or
  2. (2)
    the continuous projective geometries of Von Neumann, or
  3. (3)
    the continuous partition lattice of either Björner or Haiman,
there is some grading $\sigma$ under which $A$ is a level set.
The chief chain of $L$ (with respect to $\rho$) does not necessarily satisfy the rank modular condition with respect to the grading $\sigma$ in our construction; thus, it is not necessarily rank supersolvable in the grading $\sigma$. We do not know whether, in the situation of Theorem 1.2, it is possible to construct a grading $\sigma$ where both $A$ is a level set and also rank modularity of the chief chain is preserved.
A recent paper of Engel, Mitsis, Pelekis and Reiher [6] considers a somewhat different extension of discrete results on antichains to a continuous and/or analytic setting.
This paper is organized as follows. In Section 2, we define $R$-gradings, and review the details of the construction of Von Neumann and Björner. In Section 3, we extend facts about rank modularity and supersolvability from the discrete case to the current setting. In Section 4, we prove (an extension of) Theorem 1.2.
Acknowledgements
We’d like to thank Andrej Bauer, Sándor Radeleczki, and Alex Simpson for helpful comments.
2. Definitions and background
Let $R$ be a subset of $\mathbb{R}$, or more generally of the extended reals $\infty$$\infty$]. An $R$-rank function or $R$-grading for the poset $L$ is a function $L\to R$ that restricts to an order isomorphism on every maximal chain of $L$. Thus, a finite poset is graded (in the usual sense) if and only if it has an $[n]$-grading.
Example 2.1.
Given a complete measure space $(X,\mu)$, we take the measurable Boolean lattice $\mathcal{B}(X,\mu)$ to be the set of measurable sets up to the identification of pairs of sets whose symmetric difference has measure zero. If $\mu$ is a finite measure with no points of positive measure, then $\mu:\mathcal{B}(X,\mu)\to[0,\mu(X)]$ is a $0,\mu(X)]-$grading. We notice also, on the other hand, that the rank function on a finite Boolean lattice is exactly the counting measure on the underlying set.
The measurable Boolean lattice on the interval $[0,1]$ with Lebesgue measure is a natural continuous analogue of the Boolean lattice on $n$ points.
Our other examples come from a construction originally due to Von Neumann for modular lattices [19, 20], and extended by Björner to semimodular lattices in [2], see also [4]. We review this construction. There are two basic ingredients: a direct limit, and a metric completion.
For the direct limit, we start with a graded lattice $L_{n}$ of rank $n$ for each positive integer $n$. It will be convenient to require every $L_{n}$ to be semimodular. We renormalize the standard rank function on $L_{n}$ by dividing by $n$: thus, the top element $\hat{1}$ has renormalized rank $1$, an element in a middle rank has renormalized rank of roughly $1/2$, etc.
Whenever $k$ properly divides $n$, we consider a lattice embedding $\varphi^{k,n}:L_{k}\to L_{n}$ that respects the renormalized rank. We require that, when $k$ divides $m$ and $m$ divides $n$, it holds that $\varphi^{k,n}$ agrees with the composition of $\varphi^{k,m}$ and $\varphi^{m,n}.$ Now we take the directed limit $L_{(\infty)}$ of this system of lattices and embeddings over the positive integers with the divisibility order. In other words, we take a union over all $L_{n}$, and then identify earlier $L_{k}$’s with sublattices in later $L_{n}$’s. The renormalized rank function carries through to give a $([0,1]\cap\mathbb{Q})$-grading on $L_{(\infty)}$.
We now take a metric completion to get a $[0,1]$-graded lattice. We need to have a metric for this purpose. It has long been known that in a semimodular lattice with rank function $r$, the function $d(x,y)=2r(x\vee y)-r(x)-r(y)$ is a metric [14]. Indeed, we may view this as the length of the “up-down path” from $x$ up to $x\vee y$, and from $x\vee y$ back down to $y$. If every $L_{n}$ is semimodular, then these metrics pass to a metric on $L_{(\infty)}$. Now $L_{\infty}$ is the metric completion of $L_{(\infty)}$ under this metric. We recover a rank function by letting $\rho(x)=d(\hat{0},x)$; by definition of the distance function, this $\rho$ extends the renormalized rank function on $L_{(\infty)}$.
Example 2.2.
Von Neumann defines a continuous projective geometry over a field $\mathbb{F}$ [19, Section 5]; see also [1, Chapter X.5]. In the language above, this may be described as follows. Let $L_{n}$ be the lattice of all subspaces of the vector space $\mathbb{F}^{n}$. When $n$ is a multiple of $k$, view $\mathbb{F}^{n}$ as $\mathbb{F}^{k}\oplus\mathbb{F}^{k}\oplus\cdots\oplus\mathbb{F}^{k}$. Now define $\varphi$ by mapping a subspace $W$ of $\mathbb{F}^{k}$ to $W\oplus\cdots\oplus W$. It is easy to see that $\varphi$ preserves renormalized rank, and is a lattice mapping. Every element of the metric completion is rank modular with respect to $\rho$.
Example 2.3.
The measurable Boolean lattice on $[0,1]$ may be built from this machinery, as follows [20, Section 8]. Let $L_{n}$ be the Boolean lattice of all subsets of $[n]$. When $n$ is a multiple of $k$, view $[n]$ as $k]\times[n/k]$. Define $\varphi$ by mapping $S\subseteq[k]$ to $S\times[n/k]$, and take the directed limit and metric completion.
The identification with subsets of $[0,1]$ comes by identifying $i\in[n]$ with the interval $(\frac{i-1}{n},\frac{i}{n}]$. Since the Lebesgue measure of a union of $m$ intervals of length $\frac{1}{n}$ is $\frac{m}{n}$, the renormalized rank of $S\subseteq[n]$ agrees with the Lebesgue measure. Now $L_{(\infty)}$ consists of all sets that may be written as unions of half-open intervals with rational endpoints. The metric completion gives us unions of such sets, recovering (up to measure zero) all Borel sets.
Example 2.4.
Björner [2] builds a measurable partition lattice using a similar approach. He identifies elements of an underlying $(n+1)$-element set with the set of intervals
$\left\{\{0\},(0,\frac{1}{n}],(\frac{1}{n},\frac{2}{n},\dots,(\frac{n-1}{n},1]\right\}.$
Then when $n$ is a multiple of $k$, we identify each interval of length $\frac{1}{k}$ with $\frac{n}{k}$ intervals of length $\frac{1}{n}$ in a similar manner as in Example 2.3. Rank supersolvability of every finite partition lattice yields rank supersolvability of the limit object [2, Theorem 2].
We remark that while one might call $[0,1]$-graded or $\mathbb{R}$-graded posets “continuously graded”, they do not appear to have any relationship to speak of with the continuous posets of [10] that are studied in computability theory. Indeed, while the measurable Boolean lattice on $[0,1]$ is a main object for our purposes here, it already fails to have interesting way-below relationships. When every element of the lattice $L$ is rank modular, the $R$-grading gives $L$ the structure of a metric lattice in the sense of [1].
Haiman gives explicit descriptions of Björner’s measurable partition lattice as well as of related lattices [12]. Haiman’s lattices are again rank supersolvable [12, Lemma 2.15]. A recent preprint of Contreras Mantilla and Sinclair [5] discusses the Von Neumann and Björner construction from a model theory perspective. See also the paper [18] of Vershik and Yakubovich, and the survey article of Björner [3].
3. Rank modularity and rank supersolvable lattices
3.1. Continuity
We begin with some basic results about rank supersolvable lattices, extending the well-known theory of finite supersolvable lattices [9, 13, 16, 17].
Throughout this section, let $R$ be a subset of $\mathbb{R}$, or more generally of the extended reals $\infty$$\infty$]. Let $L$ be an $R$-graded lattice. Let $\mathbf{m}=\left\{m_{\lambda}:\lambda\in R\right\}$ be a maximal chain in $L$ with $\rho(m_{\lambda})=\lambda$, which we will take to be the chief chain of $L$ when it is rank supersolvable. We place the standard topology of $R\subseteq[-\infty,\infty]$ on each maximal chain.
We start with an easy but useful consequence of rank modularity, not requiring rank supersolvability.
Lemma 3.1.
If $w<z$ and $m^{\prime}<m$ are elements of $L$, where $m$ and $m^{\prime}$ are rank modular, then
$\rho(m\wedge z)+\rho(m\vee z)-\left(\rho(m\wedge w)+\rho(m\vee w)\right)$ $=\rho(z)-\rho(w),\text{ and}$
$\rho(m\wedge z)+\rho(m\vee z)-\left(\rho(m^{\prime}\wedge z)+\rho(m^{\prime}\vee z)\right)$ $=\rho(m)-\rho(m^{\prime}).$
Proof.
Compute! We have $\rho(m\wedge z)+\rho(m\vee z)=\rho(m)+\rho(z)$, and similarly $\rho(m\wedge w)+\rho(m\vee w)=\rho(m)+\rho(w)$. The $\rho(m)$ terms cancel, giving the first equality. The second equality is entirely similar. ∎
As in the finite case, the difference in ranks $\rho(z)-\rho(w)$ may be viewed as giving the height of the interval $[w,z]$. We recover from Lemma 3.1 several bounds on such differences.
Corollary 3.2.
If $w<z$ and $m^{\prime}<m$ are elements of $L$, where $m$ and $m^{\prime}$ are rank modular, then the following inequalities hold
$\rho(m\wedge z)-\rho(m\wedge w)\leq\rho(z)-\rho(w),\qquad$ $\rho(m\vee z)-\rho(m\vee w)\leq\rho(z)-\rho(w)$
$\rho(m\wedge z)-\rho(m^{\prime}\wedge z)\leq\rho(m)-\rho(m^{\prime}),\qquad$ $\rho(m\vee z)-\rho(m^{\prime}\vee z)\leq\rho(m)-\rho(m^{\prime}).$
Proof.
The left-hand sides of the inequalities are all nonnegative by order considerations. The left-hand sides of the first row sum to the right hand side(s) of the first row; similarly for the second row. ∎
We immediately recover two continuity properties.
Corollary 3.3.
Let $\mathbf{c}=\left\{c_{\kappa}:\kappa\in R\right\}$ be a maximal chain in $L$, with $\rho(c_{\kappa})=\kappa$. If $m$ is a rank modular element of $L$, then the following maps are Lipschitz (hence continuous at every point in $R\cap\mathbb{R}$):
[…]
Corollary 3.4.
If $L$ is rank supersolvable with chief chain $\mathbf{m}$, and $z$ is an element of $L$, then the following maps are Lipschitz (hence continuous at every point in $R\cap\mathbb{R}$):
[…]
An alternate statement of Corollary 3.3 is that the map formed by composing the three maps $\rho|_{\mathbf{c}}^{-1}$, meet with $m$, and $\rho$ is Lipschitz (hence continuous) as a map $R\to R$; similarly with Corollary 3.4.
Remark 3.5.
In the case where $L$ is semimodular, so that $d(x,y)=2\rho(x\vee y)-\rho(x)-\rho(y)$ is a metric, then it is a general fact that the join operation is Lipschitz in this metric (or a product of this metric with itself) over the entire lattice [2]. The meet operation, on the other hand, is not continuous even in the partition lattice. We can extend Corollary 3.3 to show that meet with a rank modular $m$ in a semimodular lattice is Lipschitz in the metric, since
$d(m\wedge x,m\wedge y)$ $=2\rho((m\wedge x)\vee(m\wedge y))-\rho(m\wedge x)-\rho(m\wedge y)$
$\leq 2\rho(m\wedge(x\vee y))-\rho(m\wedge x)-\rho(m\wedge y)$
$=2\rho(x\vee y)-\rho(x)-\rho(y)-2\rho(m\vee x\vee y)$
$\leq d(x,y).$
It is likely that there are more general continuity statements, but since these are incidental to our main purpose, we do not develop this idea further.
Although infinite lattices may not be bounded, this is easy to control. The following is immediate from definition.
Lemma 3.6.
The lattice $L$ has a (unique) maximum element $\hat{1}$ if and only if $R$ has a maximum element. Similarly for minimum elements.
Notice that if $R=[-\infty,\infty]$ and $\hat{1}$ is a join irreducible or $\hat{0}$ is a meet irreducible, then the rank modular relation involves an indeterminate form only when we compare the minimum $\hat{0}$ and maximum $\hat{1}$ elements, which causes no confusion.
Lemma 3.7.
If $L$ (and $R$) has no maximum element, then we can extend $L$ to a $(R\mathop{\cup}\sup R)$-graded lattice by adding a maximum element $\hat{1}$ to $L$ and taking $\rho(\hat{1})=\sup R$. If $L$ is rank supersolvable with chief chain $\mathbf{m}$, then $L\cup\{\hat{1}\}$ is rank supersolvable with chief chain $\mathbf{m}\cup\{\hat{1}\}$.
Similarly for minimum elements.
Proof.
It is immediate that the extension of $\rho$ is a grading. Since $\hat{1}\vee m=\hat{1}$ and $\hat{1}\wedge m=m$, the element $\hat{1}$ is rank modular. ∎
Corollaries 3.3 and 3.4 give continuity of certain maps at every real point of $R$. If $R$ contains $\infty$ or -$\infty$, then additional care is required. Recall from a first topology course that, since the extended reals are homeomorphic to [1,1][-1,1], a map $\varphi:[-\infty,\infty]\to[-\infty,\infty]$ is continuous if and only if it respects convergence of sequences. As an immediate consequence, we obtain the following.
Lemma 3.8.
An increasing map $\varphi:[-\infty,\infty]\to[-\infty,\infty]$ is continuous at $\infty$ (or respectively at -$\infty$) if and only if $\sup\left\{\varphi(\alpha):\alpha<\infty\right\}=\varphi(\infty)$ (or respectively, if $\inf\left\{\varphi(\alpha):\alpha>-\infty\right\}=\varphi(-\infty)$).
We remark that there are supersolvable (indeed, modular) lattices where the maps of Corollaries 3.3 and 3.4 fail to be continuous at $\pm\infty$.
Example 3.9.
Consider the product lattice $\mathbb{R}\times\mathbb{R}$ with $\mathbb{R}$-grading $\rho(a,b)=a+b$, and let $L=\mathbb{R}\times\mathbb{R}\cup\{\hat{0},\hat{1}\}$. Then the join (meet) is obtained by taking entry-wise maximums (minimums), and every element of $L$ is rank modular. Now $\mathbf{c}=\left\{(0,b):b\in\mathbb{R}\right\}\cup\{\hat{0},\hat{1}\}$ is a maximal chain where $\sup\left\{(0,b)\wedge(1,0):b\in\mathbb{R}\right\}=(0,0)$, while $\hat{1}\wedge(1,0)=(1,0)$. In particular, the map $b\mapsto\rho((0,b)\wedge(1,0))$ fails to be continuous at $\infty$. An entirely similar argument shows that the map $b\mapsto\rho((0,b)\vee(-1,0))$ fails to be continuous at -$\infty$.
Example 3.10.
Consider the lattice of all bounded Lebesgue-measurable sets in $\mathbb{R}$ (up to measure zero), with Lebesgue measure as a $0,\infty)-$grading. By modularity of measures, every element is rank modular. Then the intervals $m_{\lambda}=[-\lambda,\lambda]$ form a chief chain. Since every bounded measurable set is contained some $m_{\lambda}$, we have that $\sup_{\lambda}\left\{m_{\lambda}\wedge y\right\}=y$. Thus, the first map of Corollary 3.4 is continuous.
On the other hand, the intervals $c_{\kappa}=[1,1+\kappa]$ also form a maximal chain, and $\sup_{\kappa}\left\{c_{\kappa}\wedge[-1,1]\right\}=\{1\}\subsetneq[-1,1]$. Thus, the first map of Corollary 3.3 fails to be continuous at $\infty$.
3.2. Real-graded vs discrete-graded supersolvable lattices
For a finite lattice, there are several conditions that are equivalent to rank supersolvability. All are based on chains of elements satisfying some modularity condition. Rank functions in the non-discrete case are less rigid than those in the finite case, and that an element that is rank modular with respect to one rank function does not generally imply that it will be rank modular with respect to another.
Example 3.11.
If $\rho(x)$ is a $[0,1]$-grading on $L$, then $\left(\rho(x)\right)^{2}$ is another $[0,1]$-grading. If $m$ is rank modular with respect to $\rho$, however, then (by the failure of the Freshman’s Dream identity in characteristic 0) it typically will not be rank modular with respect to $\rho^{2}$.
Example 3.12.
If $f:\mathbb{R}\to\mathbb{R}$ is a strictly increasing function, and $\rho$ is an $R$-grading, then $f\circ\rho$ is an $f(R)$-grading. If $f$ is affine linear, then $m$ is rank modular with respect to $\rho$ if and only if it is rank modular with respect to $f\circ\rho$. If $f$ is not affine linear, then it typically does not preserve rank modularity of elements.
Going from rank modularity to other modularity conditions is easier. The next lemma gives that a rank modular element is “left modular”.
Lemma 3.13.
If $m$ is rank modular in $L$, then for any $w<z$ in $L$, it holds that
$(w\vee m)\wedge z=w\vee(m\wedge z).$
Proof.
Apply exactly the same argument as in [9, Lemma 4.1]. ∎
Thus, if $m$ is rank modular, then we may write $w\vee m\wedge z$ without parentheses.
A maximal chain consisting of rank modular elements gives somewhat stronger lattice conditions:
Lemma 3.14.
Let $L$ be rank supersolvable with chief chain $\mathbf{m}$. For $z<w\in L$ and $m_{\lambda}<m_{\kappa}\in\mathbf{m}$, it holds that
$(z\vee m_{\lambda})\wedge w$ $=z\vee(m_{\lambda}\wedge w),\text{ and}$
$(m_{\lambda}\vee z)\wedge m_{\kappa}$ $=m_{\lambda}\vee(z\wedge m_{\kappa}).$
Proof.
The first identity is a special case of Lemma 3.13.
For the second identity, we notice that since $(m_{\lambda}\vee z)\wedge m_{\kappa}\geq m_{\lambda}\vee(z\wedge m_{\kappa}),$ it suffices to show that the ranks agree. Now, since $m_{\lambda}\vee z\vee m_{\kappa}=z\vee m_{\kappa}$ and $m_{\lambda}\wedge z\wedge m_{\kappa}=m_{\lambda}\wedge z$, we get that
$\rho((m_{\lambda}\vee z)\wedge m_{\kappa})$ $=\rho(m_{\lambda}\vee z)+\kappa-\rho(z\vee m_{\kappa})$
$\rho(m_{\lambda}\vee(z\wedge m_{\kappa}))$ $=\lambda+\rho(m_{\kappa}\wedge z)-\rho(m_{\lambda}\wedge z).$
Now Lemma 3.1 gives the difference of these ranks to be zero, as desired. ∎
We observe in passing that since the sublattice generated by a finite modular chain and another finite chain forms a distributive lattice [9, 16], the sublattice of $L$ generated by (finite) combinations of $\mathbf{m}$ and another chain $\mathbf{c}$ is distributive.
In the finite case, the converse of Lemma 3.14 also holds true [9, Theorems 1.2 and 1.4]. The difficulty in extending this converse to the non-discrete case is in constructing a nice enough rank function.
Question 3.15.
Given an $R$-graded lattice $L$ and a chain $\mathbf{m}$ satisfying the conclusion of Lemma 3.14, is it possible to find a rank function under which all elements of $\mathbf{m}$ are rank modular?
We close this section by extending one more result from finite lattices to $R$-graded lattices: rank modular elements project into intervals nicely.
Lemma 3.16.
If $m$ is a rank modular element of $L$, then $w\vee m\wedge z$ is a rank modular element of the interval $[w,z]$ for any $w<z$ in $L$.
Proof.
It suffices by duality and iteration to show that $m\wedge z$ is rank modular in the interval $\hat{0},z]$. For $w<z$, we desire to have $\rho(w\vee m\wedge z)=\rho(w)+\rho(m\wedge z)-\rho(m\wedge w)$. But the rank modular identity with $m$ and $w\vee m\wedge z$ gives us that
$\rho(m\vee(w\vee m\wedge z))+\rho(m\wedge(w\vee m\wedge z))-\rho(m)=\rho(w\vee m\wedge z).$
Noticing that $m\vee(w\vee m\wedge z)=m\vee w$ and $m\wedge(w\vee m\wedge z)=m\wedge z$, and expanding $\rho(m\vee w)$ as $\rho(m)+\rho(w)-\rho(w\wedge m)$ now yields the desired. ∎
4. Proof of Theorem 1.2
Throughout this section, let $L$ be a rank supersolvable lattice with a given $R$-grading $\rho$ over some interval $R\subseteq\mathbb{R}$. We may assume without loss of generality by Lemma 3.7 that $R$ is compact (passing to the extended reals if necessary), and so that $L$ has a minimum element $\hat{0}$ and maximum element $\hat{1}$. Let $\mathbf{m}=\left\{m_{\lambda}:\lambda\in R\right\}$ be the chief chain guaranteed by rank supersolvability, where $\rho(m_{\lambda})=\lambda$. Let $A$ be an antichain cutset.
We will prove the following generalization of Theorem 1.2.
Theorem 4.1.
Let $R\subseteq[-\infty,\infty]$ be an interval. Let $L$ be a rank supersolvable lattice with respect to the $R$-grading $\rho$, with chief chain $\mathbf{m}$. Suppose further that the following conditions are met for every maximal chain $\mathbf{c}$, $\lambda_{0}\in R$, and $z\in L$:
$\sup\big\{m_{\lambda_{0}}\wedge c$ $:c\in\mathbf{c},\rho(c)<\infty\big\}=m_{\lambda_{0}}\quad$ $\quad\inf\big\{m_{\lambda_{0}}\vee c$ $:c\in\mathbf{c},\rho(c)>-\infty\big\}=m_{\lambda_{0}}$
$\sup\big\{m_{\lambda}\wedge z$ $:\lambda\in R,\lambda<\infty\big\}=z$ $\inf\big\{m_{\lambda}\vee z$ $:\lambda\in R,\lambda>-\infty\big\}=z.$
If $A$ is an antichain cutset for $L$, then there is some grading $\sigma$ under which $A$ is a level set.
We notice that if $R$ is bounded from above (in $\mathbb{R}$), then the supremum conditions are automatic; similarly if $R$ is bounded from below with the infimum conditions. Indeed, in light of Lemma 3.8 and the discussion surrounding, we see that the additional conditions give exactly that the maps of Corollaries 3.3 and 3.4 are continuous also at $\infty$ and/or at -$\infty$ (if present in $R$). We summarize with a lemma:
Lemma 4.2.
Under the conditions of Theorem 4.1, writing $\mathbf{c}=\left\{c_{\kappa}:\kappa\in R\right\}$, the following maps $R\to R$ are continuous.
$\lambda\mapsto\rho(m_{\lambda}\wedge z)\qquad$ $\lambda\mapsto\rho(m_{\lambda}\vee z)$
$\kappa\mapsto\rho(m_{\lambda_{0}}\wedge c_{\kappa})\qquad$ $\kappa\mapsto\rho(m_{\lambda_{0}}\vee c_{\kappa})$
Remark 4.3.
Examination of Examples 3.9 and 3.10 demonstrates that continuity at $\pm\infty$ is not automatic. Since the lattice of Example 3.9 is isomorphic to a space-time causality order, the conclusion of Theorem 4.1 nonetheless holds by [7]. We do not know whether the antichain cutsets of the lattice in Example 3.10 are level sets in some $0,\infty)-$grading.
We now construct the new rank function $\sigma$ for Theorem 4.1. Consider the good chain through $z$ given by $\left\{z\wedge m_{\lambda}\,:\,\lambda\in R\right\}\cup\left\{z\vee m_{\lambda}\,:\,\lambda\in R\right\}$. We will show that the good chain through $z$ is a maximal chain, hence it intersects with $A$ at a unique point $\alpha(z)$. We define
$\sigma(z):=\rho(z)-\rho(\alpha(z)).$
[…]
First, in order for $\sigma$ to be well-defined, we need to check that the good chain is maximal.
Lemma 4.4.
The good chain through $z$ is a maximal chain.
Proof.
It suffices to show that elements of the good chain achieve every rank in $R$.
It follows from Lemma 4.2 that $\rho(\left\{z\wedge m_{\lambda}\,:\,\lambda\in R\right\})$ is the continuous image of a connected set having upper bound $\rho(z)$; similarly for $\rho(\left\{z\vee m_{\lambda}\,:\,\lambda\in R\right\})$. As these share the point $\rho(z)$, the ranks of a good chain form a connected subset of $R$. Since $\hat{0}=\hat{0}\wedge z$ and $\hat{1}=\hat{1}\vee z$, the desired follows. ∎
An entirely similar proof gives:
Lemma 4.5.
If $\mathbf{c}$ is a maximal chain and $m$ is on the given chief chain $\mathbf{m}$ of $L$ , then the chain $\{m\wedge c:c\in\mathbf{c}\}\cup\{m\vee c:c\in\mathbf{c}\}$ is a maximal chain.
Remark 4.6.
By Corollary 3.3, if $L$ is $R$-graded for $R$ a bounded subset of $\mathbb{R}$, then we may relax the requirement that $m$ lie on the given chief chain to simple rank modularity.
We will mostly focus on elements $z$ above the antichain cutset $A$; elements below the antichain cutset may be dealt with in a completely symmetrical manner.
We start with an easy observation:
Lemma 4.7.
If $w<z$ are above or on the antichain cutset, so that $\alpha(w)=m_{\lambda}\wedge w$ and $\alpha(z)=m_{\kappa}\wedge z$, then either $\lambda>\kappa$ or else $\alpha(w)=\alpha(z)$.
Proof.
If $\lambda\leq\kappa$, then $m_{\lambda}\wedge w\leq m_{\kappa}\wedge z$. Since both lie on the antichain $A$, equality follows. ∎
See the illustration in Figure 4.1. We can now obtain the increasing property:
Lemma 4.8.
The map $\sigma$ is strictly increasing.
Proof.
We need to show that if $w<z$, then $\sigma(w)<\sigma(z)$.
We first check the case where $w$ and $z$ are both above or on $A$. We write $\alpha(w)=m_{\lambda}\wedge w$ and $\alpha(z)=m_{\kappa}\wedge z$. If $\alpha(w)=\alpha(z)$, then the result is immediate since $\rho$ is strictly increasing. Thus, we may assume by Lemma 4.7 that $\lambda>\kappa$. Now
$\sigma(z)-\sigma(w)$ $=\left(\rho(z)-\rho(w)\right)-\left(\rho(m_{\kappa}\wedge z)-\rho(m_{\lambda}\wedge w)\right)$
$>\left(\rho(z)-\rho(w)\right)-\left(\rho(m_{\kappa}\wedge z)-\rho(m_{\kappa}\wedge w)\right)$
$\geq 0.$
where the strict inequality is because $m_{\kappa}\wedge w=\alpha(w)\wedge\alpha(z)<\alpha(w)$, and the weak inequality is by Corollary 3.2.
Finally, if both are below or on $A$, then the result follows by duality. If $w$ is below $A$ and $z$ is above, then there is by definition an element $x$ of $A$ on a maximal chain between them, and then $\sigma(w)<\sigma(x)<\sigma(z)$. ∎
[…]
Figure 4.1. Projection to $A$.
Lemma 4.9.
If $\mathbf{c}$ is a maximal chain, then $\hat{\alpha}=\rho\circ\alpha\circ\rho|_{\mathbf{c}}^{-1}$ is a continuous map $R\to\mathbb{R}$.
That is, the rank of $\alpha(z)$ varies continuously over the chain $\mathbf{c}$.
Proof.
We prove that $\hat{\alpha}$ is continuous from the right at a point $w$ that is above or on the antichain cutset $A$. Continuity from the left above $A$ is entirely similar, and points below or on the antichain cutset follow from duality.
By Lemma 4.2 and the fact that the continuous preimage of a point is closed, there is a minimal $\lambda$ with $m_{\lambda}\wedge w=\alpha(w)$. It follows that $m_{\lambda}\geq\alpha(w)$. If $m_{\lambda}=\alpha(w)$, then $m_{\lambda}=\alpha(z)$ also for every $z>w$, and a constant function is right-continuous.
Thus, we may assume that $m_{\lambda}>\alpha(w)$. By Lemmas 4.2 and 4.4 and minimality of $\lambda$, we can choose an $m_{\kappa}$ that is above the antichain cutset, below $m_{\lambda}$, and so that $\rho(m_{\lambda}\wedge w)-\rho(m_{\kappa}\wedge w)<\epsilon/2$. By Lemma 4.5, the chain $\left\{m_{\kappa}\wedge c:c\in\mathbf{c}\right\}$ is maximal on the interval $\hat{0},m_{\kappa}]$, hence there is some $z_{0}>w$ so that $m_{\kappa}\wedge z_{0}=\alpha(z_{0})$. Now, since $m_{\kappa}\wedge z\leq\alpha(z_{0})$ and $m_{\lambda}\wedge z\geq\alpha(w)$, every $z$ between $w$ and $z_{0}$ has $\alpha(z)$ between $m_{\kappa}\wedge z$ and $m_{\lambda}\wedge z$. In particular, if $z\in\mathbf{c}$ is chosen so $\rho(m_{\lambda}\wedge z)<\min(\rho(m_{\lambda}\wedge w)+\epsilon/2,\rho(m_{\lambda}\wedge z_{0}))$, then $\alpha(z)$ and $\alpha(w)$ are both on an interval $m_{\kappa}\wedge w,m_{\lambda}\wedge z]$. Since $\rho(m_{\lambda}\wedge z)-\rho(m_{\kappa}\wedge w)<\epsilon$, the result now follows. ∎
Corollary 4.10.
[…]
Proof.
[…]
Theorem 1.2 now follows from Lemma 4.8 and Corollary 4.10.
Example 4.11.
The chief chain may not be rank modular with respect to the new rank function $\sigma$. Consider the measurable Boolean lattice $\mathcal{B}$ on $[0,2]$ with Lesbegue measure $\mu$. Let $\nu(U)$ be the integral over $U$ of the function that is $1$ on $[0,1]$ and $2$ on $[1,2]$, and let $A$ be the family of all sets of measure $1$ with respect to $\nu$, an antichain cutset in $\mathcal{B}$. Since $\mathcal{B}$ is rank modular, any maximal chain is a chief chain. We take the chief chain consisting of the intervals $[0,t]$ for $0\leq t\leq 2$.
But now $\sigma([0,t])=\mu([0,t])-\mu([0,1])=t-1$, while $\sigma([1,2])=\mu([1,2])-\mu([1,3/2])=1/2$. Now $\sigma([0,1])+\sigma([1,2])\neq\sigma([0,2])+\sigma(\emptyset)$. Thus, the element $[0,1]$ is not rank modular with respect to $\sigma$!